1.If the market price is $10, what is the firm’s s

1.If the market price is $10, what is the firm’s short-run economic profit?

a. $9 b. $15 c. $30 d. $50

2.If the market price is $6, what is the firm’s short-run economic profit?

a. $0 b. $12 c. $15 d. $18

3.If the market price is $10, what is the firm’s total cost?

a. $15 b. $30 c. $35 d. $50

4.If the market price is $10, what is the firm’s total revenue?

a. $15 b. $30 c. $35 d. $50

5.The firm will earn zero economic profit if the market price is

a. $0. b. $6. c. $7. d. $10.

6.The firm will earn positive economic profit if the market price is

a. positive.

b. $6.
c. above $6

d. There is no price at which the firm earns positive economic profits.
Answer #1

1.

Correct Answer:

B. $15

In competitive market,

P = MR = MC = $10

Q = 5

Economic profit = total revenue – total cost = 10*5 – 7*5 = $15

2.

Correct Answer:

A. $0

P = MR = MC = $6

Economic profit = total revenue – total cost = 6*3 – 6*3 = 0

3.

Correct Answer:

C. $35

When price is $10.

Total cost = Average cost * quantity = 7*5 =$35

4.

Correct Answer:

D. $50

When price is $10.

Total revenue= Average price * quantity = 10*5 =$50

5.

Correct Answer:

B. $6

At price = $6

Economic profit = total revenue – total cost = 6*3 – 6*3 = 0

6.

Correct Answer:

C. above $6

Above the price of $6, firm will earn positive economic profit.